Write an equation for each number:
1. A rectangle and an equilateral triangle have the same perimeter. The length of the rectangle is 3 cm more than the width. Each side of the triangle is 10 cm. Find the length and width of the rectangle.
2. The Jones family is taking a trip from Seattle, WA to San
Diego, CA over the course of 3 days. They plan on traveling 200 miles more on
the second day than they will on the first day. They will travel 75 miles less on the
third day than they will on the first day. To total distance of the trip is 1058 miles.
How many miles will they travel each day?

Answer by Boreal(15235)

The triangle has 3 equal sides of 10 cm for a perimeter of 30 cm
The rectangle has 2 sides of x and 2 sides of x+3 for a perimeter of 4x+6=30 cm
So x, the length of the short side, is 6 cm and the long side is 9 cm.
-
x miles first day
x+200 on second day
x-75 on third day
add them: 3x+125=1058
3x=933
x=311 miles first day
x+200=511 miles second day
x-75=236 miles third day
They add to 1058 miles